Jubilee Feed and Supply Company buys on term, of 1/10 net 30, but it has not been taking discounts and has actually been paying in 60 rather than 30 days. Assume that the accounts payable are recorded at full costs, not net of discounts. Jubilees statement of financial position follows: (thousands of shillings). ASSETS Cash Accounts receivable Inventory Total Current Assets Kshs. LIABILITIES EQUITY Kshs. Kshs.5,000 500 500 12,500 Total Current Liabilities Ksh.6.000 Kshs.500 Accounts payable 4,500 Notes payable Accruals Property, plant and Equipment Total Assets Z.500 Long-term debt 20,000 Shareholders equity 1,500 12.500 Total Liabilities & Equity 20,000 Now, Jubilees supplies are threatening to stop shipment unless the company begins making prompt payments (that is, paying in 30 days or less). The firm can borrow on a 1-year note (call this a current liability) fron its bank at a rate of 15%, with a 20% compensating balance required. Jubilees kshs, 500,000of cash is needed for transactions; it cannot be used as part of the compensating balance) Required: a) How large would the accounts payable balance be if Jubilee takes discounts? If it does not take discounts and pays in 30 days (10 marks) How large must the bank loan be if Jubilee takes discounts? If Jubilee doesnt take discounts? b) (5 marks) c) Assume that Jubilee foregoes the discount and borrows the amount needed to become current on its payables. Constant a Proforma statement of financial position based on this decision (5 marks) d) Now assume that the Kshs.5,000,000 shows on the statement of financial position is recorded net of discounts. How much would Jubilee have to pay its suppliers to reduce its accounts payable to Kshs2,500,000? If Jubilees tax rate is 30%,what is the effect on its net income due to the lost discount when it reduces its accounts payable to Kshs.2,500,000?How much would Jubilee have to borrow? (5 marks)
Answer #1

a)

Account balance if Jubilee takes discount:

Account Balance = 5,000,000 * 99% = 4,950,000

Account balance if Jubilee does not takes discount:

Account Balance = 5,000,000 * 100% = 5,000,000

Note : It has been mentioned that values has been recorded on gross value.

b)

Bank loan amount if Jubilee takes discount:

Bank loan = Loan amount + 20% of Loan amount

Let us assume Loan amount = x

Then

X=4,950,000 + X * 20%

X = 6,187,500

Loan amount = 6,187,500

Bank loan amount if Jubilee does not take discount:

Bank loan = Loan amount + 20% of Loan amount

Let us assume Loan amount = x

Then

X=5,000,000 + X * 20%

X = 6,250,000

Loan amount = 6,250,000

c)

Financial position when Jubilee forgoes discount:

Balance sheet (Kashs in '000)

Assets

Kashs

Liability and Equity

Kashs

Cash

500

Notes Payable

6,750

Account receivable

4,500

Accruals

500

Overdraft available

1,250

Total current liability

7,250

Inventory

7,500

Total current assets

13,750

Long term debt

1,500

Property Plant and Equipment

7,500

Shareholder equity

12,500

Total Assets

21,250

total equity and liability

21,250

Note1: Total notes payable has been shown with compensatory balance.

Note2: Overdraft available has been shown in the assets side of 1,250,000

d)

Account payable net of discount is 5,000,000

Gross account payable = 5,000,000/.99 = 5,050,505

To reduce account payable up to 2,500,000 Jubilee must pay 2,550,505 (5,050,505 – 2,500,000)

Let us assume Loan amount = x

Then

X=2,550,505 + X * 20%

X = 3,188,131.25

Loan amount = 3,188,131.25

Amount of lost discount = 5,050,505-5,000,000 = 50,505

Decrease in tax liability due to discount lost = 50,505 * 30% = 15,151.50

Discount lost = profit lost = 50,505.

Note: It has been assumed that discount has not been availed and account payable is adjusted to undiscounted value.